OpenStudy (needhelpp101):

Solve each quadratic equation. Problem: (2x – 1)(x + 7) = 0 Answer: x= ½, -7 Can you please show me how to solve it? Show your work! Work it out!

OpenStudy (agent0smith):

Set each factor equal to zero and solve. Here's an example: (4x + 5)(x - 12) = 0 4x + 5 = 0 or x - 12 = 0 x = -5/4 or x = 12